OpenStudy (anonymous):

Witch Fraction Is Equivalent To 5 3/4? A.19/4 B.15/4 C.23/4 D.8/4

OpenStudy (alexandervonhumboldt2):

\[a \frac{ b }{ c }=\frac{ a*c+b }{ c }\]

OpenStudy (alexandervonhumboldt2):

do the same with your fration

OpenStudy (johan14th):

What number is divisible by 4 5 times that it will leave you with a remainder of 3 after you divide?

OpenStudy (anonymous):

C

OpenStudy (johan14th):

yes 23/4 =5 and 3 is left over which equals to 5 (3/4) :)